# Effects of particle angularity on granular self-organization

### v1

@@ Abstract
Recent studies of two-dimensional poly-disperse disc systems revealed a coordinated self-organisation of cell stresses and shapes, with certain distributions collapsing onto a master form for many processes, size distributions, friction coefficients, and cell orders. Here we examine the effects of grain angularity on the indicators of self-organisation, using simulations of bi-disperse regular $N$-polygons and varying $N$ systematically. We find that: the strong correlation between local cell stresses and orientations, as well as the collapses of the conditional distributions of scaled cell stress ratios to a master Weibull form for all cell orders $k$, are independent of angularity and friction coefficient. In contrast, increasing angularity makes the collapses of the conditional distributions sensitive to changes in the friction coefficient.
@@ I Introduction
Understanding behaviour of granular matter is essential in many fields of research and technology. Macroscale models of stress transmission in such media involve upscaled constitutive properties that often presume independence of the stress field, which enables explicit solutions. The constitutive properties are sensitive to the local grain-scale structural characteristics, but these were shown recently to be strongly correlated with the local stress due to a cooperative local stress-structure self-organization Matsushima and Blumenfeld (2017, 2021); Jiang et al. (2022). Significantly, this calls into question models involving linear stress field equations. While relation between macroscopic properties and grain-scale characteristics, such as force-chains, have been studied extensively, advances on effects of dynamics on the settled structure have been limited Nguyen et al. (2009); Zhu et al. (2016a, b). Recent studies of planar (2D) disc packings have shown that the basic organizable structural elements, cells, which are the smallest voids surrounded by grains in contact, self-organize during slow dynamics in ways that transcend particular system and process details Matsushima and Blumenfeld (2014, 2017, 2021); Sun et al. (2020); Jiang et al. (2022), including remarkable steady-state features Wanjura et al. (2020); Sun et al. (2021); Myhill and Blumenfeld (2023); Wanjura et al. (2024).
These studies beg the obvious question how the self-organisation (SO) is affected by grain angularity. Not only are non-spherical grains are more ubiquitous in nature but they also restrict mobility and affect differently the stability of granular aggregates. However, because of the involved analytical and numerical treatments of such shapes, a limited body of work exists on such systems, e.g. Azéma et al. (2012); Krengel et al. (2023); Jiang et al. (2024); Hua et al. (2022). While the combined effects of grain shape and friction on bulk properties have received some attention Binaree et al. (2020, 2021); Krengel et al. (2025), the micro-structural effects on cell characteristics in 2D are not fully understood, with work so far limited to squares with rounded corners Jiang (2024).
A way to explore effects of particle angularity in 2D, while making connection with discs is by studying $N$-sided regular polygonal grains and varying $N$ systematically. The polygons approach discs in the limit $N\to\infty$. Here, we investigate systematically which of the observed indicators of SO Krengel et al. (2023); Jiang et al. (2021, 2022) survive the increasing angularity, as well as the interplay between angularity and friction.
@@ II.1 Simulation setup
We performed isotropic biaxial compression tests on systems of regular convex $N$-polygons with $N=5,6,13$, and $64$ the number of polygon edges. In our simulations, we used the hard-particle soft-contact Discrete Element Method Matuttis and Chen (2014); Krengel et al. (2023). In each simulation, 2200-2800 frictionless particles of equal shape were deposited under gravity to generate dense initial packings. To minimize crystallization, each assembly contained a mixture of 80% small and 20% large particles. Once generated, gravity was disabled and the systems were compressed uniaxially until the pressure on all four walls reached $20$kPa. Energy was then allowed to dissipate until pressure fluctuations decayed to less than $10^{-14}$ of the constant pressure. At this ’initial’ state, intergranular friction, $\mu$, was switched on and the packing was compressed biaxially in the direction of gravity, keeping the side walls at constant pressure of $20$kPa, see Fig. 1. Each combination of $N$ and $\mu$ was run five times, starting from different initial random configurations. In each simulation, we varied $\mu$ in the range $[0.01,10]$ for a total of 20 different assemblies. The pure shear, generated by the biaxial compression continued until, and beyond, the system reached a steady state of the global stress. We checked that, with one exception, which we discuss in detail below, no measurable quantity reached a steady state in any of the runs later than $8$% strain. All data were collected at the steady state between strains of $8$% and $16$%.
Figure 1: The biaxial shear setup and boundary conditions. The top wall is lowered down at a constant speed $v_{y}$ and the side walls are free to move to maintain a constant confining pressure $\sigma_{c}$.
@@ II.2 Cells and cell stresses
A cell $c$ is a polygon, whose vertices are the centroids of the overlapping areas of its surrounding grains, shown as the red shaded area in Fig. 2. These centroids are regarded as ’contact points’. The cell polygons in our simulations were generically convex, although this need not generally be the case  [^1] Cell convexity is realised when surface ’point contacts’ are defined as the middle points of the parts of the surfaces in contact.. The order of a cell, $k(\geq 3)$, is defined as the number of grains surrounding it. Our evaluation is limited to cells of order $k\leq 8$ in the absence of good statistics for the rare and short-lived higher order cells.
We also define the cell stress $\bm{\sigma}^{c}$, based on the stresses of the grains surrounding it, as follows. A grain stress is
$\bm{\sigma}^{p}=\frac{1}{A^{p}}\sum_{p^{\prime}}\bm{l}^{pp^{\prime}}\otimes\bm{f}^{p^{\prime}p},$ (1)
with $p^{\prime}$ the grains in contact with $p$, $\bm{l}^{p^{\prime}p}$ vectors connecting the centre of $p$ with the contact with $p^{\prime}$, $\bm{f}^{p^{\prime}p}$ the contact force between $p$ and $p^{\prime}$. The grain area, $A^{p}$, is the sum of the areas of its quadrons Blumenfeld (2004); Blumenfeld and Edwards (2006): $A^{pc},A^{pc^{\prime}},A^{pc"}$, which extend to its surrounding cells, respectively, $c,c^{\prime},c"$, as shown in Fig. 2. The cell stress is then defined as
$\bm{\sigma}^{c}=\frac{\sum_{p\in c}A^{pc}\bm{\sigma}^{p}}{\sum_{p\in c}{A^{pc}}},$ (2)
with $p\in c$ denoting the polygons surrounding cell $c$.
Figure 2: An example of a cell of order $k=8$: The cell is represented as a polygon, whose vertices are the centroids of the overlapping areas of its surrounding grains. $A^{pc},A^{pc^{\prime}}$, and $A^{pc"}$ are the areas of the quadrons of $p$ extending into its surrounding cells, respectively, $c,c^{\prime}$, and $c"$. The cell area is the sum of the quadrons of the grains extending into it.
@@ III Results
We find that the conditional cell order probabilities, $P(k|N)$, depend on on both the angularity and on friction (Fig. 3) – the more angular the grains the more sensitive they are to $\mu$. Interestingly, while cells of order $3$ are the most frequent for $N\leq 13$, cells of order $4$ are the most frequent when $N$ increases. As expected, increasing $\mu$ also increases the occurrence probabilities of high-order cells, but up to cell order $5$, $P(k\leq 5|N)$ decreases with $\mu$ for all $N$. Reducing $N$ increases the proportion of order-$3$ cells and more compact structures. For high $k$, $P(k\leq 5|N)$ is found to depend somewhat on $\mu$ and hardly at all on $N$. The rates of merging and breaking of cells fluctuate about the steady-state values and these fluctuations increase with both $\mu$ and $N$ (see Fig. 3 in the supplementary material (SM) SM ).
@@
Denoting the principal cells stresses $\sigma_{1}^{c}$ and $\sigma_{2}^{c}<\sigma_{1}^{c}$, $h=\sigma_{2}^{c}/\sigma_{1}^{c}$ is a measure of cell stability. The conditional probability density function (PDF) of $h$, $P(h|k)$, was found in Jiang et al. (2022) to collapse to a Weibull distribution in disc systems for all $k$ and $\mu$, when scaled by the mean $\bar{h}_{k}$, $P\left(\hat{h}\equiv{h}/{\bar{h}_{k}}|k\right)=P\left(\hat{h}\right)$ $\forall k$. Scaling, for each value of $\mu$, by $\bar{h}_{k}(\mu)$, find a good collapse for all $k$, shown in Fig. 4. Scaling by the average of $\bar{h}_{k}(\mu)$ over all values of $\mu$ yields a relatively good collapse for $N=64$, but the collapse’s quality deteriorates with angularity. These collapses are also fitted well by the Weibull form
$P\left(\hat{h}\right)=\frac{m}{\lambda}\left(\frac{\hat{h}}{\lambda}\right)^{m-1}e^{-(\bar{h}/\lambda)^{m}}\ ,$ (3)
with $m$ and $\lambda$ hardly depending on $N$ (see Fig. 4 in the SM SM ), $m=3.1\pm 0.1$ and $\lambda=0.949\pm 0.002$.
Figure 4: The collapsed form of $P(\hat{h})$ for different values of $\mu$ and $N$. The dashed lines represent Weibull-distribution fits for the different values of $\mu$.
In Fig. 5 we plot the mean cell stress ratio for each $k$, $N$, and $\mu$. While the initial evolution of the cell stress-ratio, $\sigma_{1}/\sigma_{2}$, is qualitatively different for the different combinations of $N$ and $\mu$, the steady-state values were very similar. This is further highlighted in Fig. 5 in the SM SM . The exception from the above is the most angular grains assembly, $N=5$, which we discuss in detail below and in the concluding section. Nevertheless, the convergence in almost all cases to a very similar steady-state stress ratio, observed here for the first time, supports further the idea of cooperative stress-structure self-organisation Jiang et al. (2022).
@@
In Fig. 5 in the SM SM , we see that the steady-state values of $\hat{h}$ first increase with $\mu$ and then decrease, with the decrease more pronounced with increasing angularity and almost disappearing for $N=64$, when $k=8$. We confirm earlier observations Krengel et al. (2025) that $\hat{h}$ peaks at $\mu=0.5$. This maximum is the result of two competing mechanisms, as we discuss in more detail below.
To investigate cell orientational organization, we approximate each cell as the ellipse closest to its contacts Jiang et al. (2022) (Fig. 6). We measured the ellipses’ aspect ratios, $\alpha$, and major axes orientations, $\theta^{c}$. We omit in this analysis cells of order $k=3$ for which the ellipse approximation is not useful.
@@
Figure 7 shows the evolution of $\alpha$ with strain for combinations of $N$ and $\mu$. At low friction, $\alpha$ is constant regardless of $N$. As $\mu$ increases the mean aspect ratio increases and then decrease to a steady state value. The maximum is more pronounced for large $k$, low $N$, and high $\mu$. It is hardly observable or disappears altogether for $k\leq 5$. In this regime, $\alpha$ decreases with increasing cell order. This can be understood in the context of the entropy-mechanical stability competition Sun et al. (2020), with the latter dominating at high-order cells.
@@
Figure 8: Cell aspect ratios at steady state as a function of $\mu$ for different angularity $N$.
@@
Cell organization is driven by stresses and we find that, for all $N$ and $\mu$, there is a strong correlation between cell orientations $\theta^{c}$ (see also Fig. 6 in the SM SM ) and the local major stress principal axis $\theta^{S}$. In Fig. 9 we plot the distribution of the local difference between the two orientations, $\Delta\theta\equiv\theta^{c}-\theta^{S}$, which shows clearly their tendency to align. This phenomenon, which is more pronounced at high-order cells, is as in disc systems Matsushima and Blumenfeld (2021); Jiang et al. (2022), suggesting that it is independent of grain angularity.
@@
We understand the dynamics that lead to such organisation as follows: Although the process is quasistatic, reorientation of cells requires unbalanced forces to move particles. This means that the cell stresses are momentarily asymmetric, giving rise to small torques that rotate them. The less stable the cell the larger the rotation. A measure of stability is the ratio of the mean shear stress, $\tau=$, and the trace, $p$, $\tau/p=(\sigma_{1}-\sigma_{2})F\sin{(2\Delta\theta)}/(\sigma_{1}+\sigma_{2})$. Thus, the nudging of the less stable cells towards higher stability during the process reduces $\Delta\theta$ and gives rise to the observed peak around $\Delta\theta=0$. This can also be seen in the scatter plot of cells in the plane spanned by $\Delta\theta$ and $\tau/p$, shown in Fig. 10.
Figure 9: PDF of $\Delta\theta\equiv\theta^{c}-\theta^{S}$ for different cell orders, $k$.
@@
Figure 10: A scatter plot heat map of cells in the plane spanned by $\Delta\theta$ and $\tau/p$ plane. The color denotes the relative frequency of cells with given values of $\Delta\theta$ and $\tau/p$.
Overall, the behaviour of the most angular grain systems, $N=5$, differs from the one common to discs and high-$N$ systems in several aspects, stemming from a competition between two cell rearrangement mechanism Krengel et al. (2025): At low friction, while particle rearrangement is dominated by sliding, the surrounding granular matrix is easily deformable and particle rotation is induced by normal collisions. As the inter-particle friction increases, sliding across flat surfaces becomes increasingly restricted and frictional induced rotation starts to gain importance. However, for angular particles, the tangential force barrier to induce rotation is higher than for high-$N$ particles, while at the same time larger contact interfaces dissipate more energy, so for a given value of $\mu$ contacts between angular particles are more stable than contacts between round particles. This in turn means, that higher order cells and larger cell aspect ratios are more likely for more angular grains.

### v2

# Effects of particle angularity on granular self-organisation

@@ Abstract
Recent studies of two-dimensional poly-disperse disc systems revealed a coordinated self-organisation of cell stresses and shapes, with certain distributions collapsing onto a master form for many processes, size distributions, friction coefficients, and cell orders. Here we examine the effects of particle angularity on the indicators of self-organisation, using simulations of bi-disperse regular $N$-polygons and varying $N$ systematically. We find that: the strong correlation between local cell stresses and orientations, as well as the collapses of the conditional distributions of scaled cell stress ratios to a master Weibull form for all cell orders $k$, are independent of angularity and friction coefficient. In contrast, increasing angularity makes the collapses of the conditional distributions sensitive to changes in the friction coefficient.
@@ I Introduction
Understanding behaviour of granular matter is essential in many fields of research and technology. Macroscale models of stress transmission in such media involve upscaled constitutive properties that often presume independence of the stress field, which enables explicit solutions. The constitutive properties are sensitive to the local particle-scale structural characteristics, but these were shown recently to be strongly correlated with the local stress due to a cooperative local stress-structure self-organisation [1, 2, 3]. Significantly, this calls into question models involving linear stress field equations. While relation between macroscopic properties and particle-scale characteristics, such as force-chains, have been studied extensively, advances on effects of dynamics on the settled structure have been limited [4, 5, 6]. Recent studies of planar (2D) disc packings have shown that the basic organisable structural elements, cells, which are the smallest voids surrounded by particles in contact, self-organise during slow dynamics in ways that transcend particular system and process details [7, 1, 2, 8, 3], including remarkable steady-state features [9, 10, 11, 12].
Self-organisation (SO) is a ubiquitous phenomenon, observed in many aspects of life, from natural phenomena to social behaviour [13, 14, 15, 16, 17]. Its emergence depends on a wide range of variables, some of which are continuous. This makes a general theory difficult, although some attempts appeared recently [16, 18, 19]. In the context of granular dynamics, many signatures of SO have been observed in the last few years and one question is how these are affected by particle angularity. Not only are non-spherical particles more ubiquitous in nature [20] but they also restrict mobility [21, 22, 23, 24] and affect the stability of granular aggregates [25, 26, 27, 28]. Owing to the involved analytical and numerical treatments of such shapes, a limited body of numerical work exists on micro- and meso-scale such systems, e.g. [29, 30, 31, 32, 33, 34, 35]. In contrast, the effects of inter-particle friction has received more attention [36, 37, 8, 38, 39, 40], although primarily in systems of discs and spheres. However, the effects of shape and friction on the behaviour of particles are not independent of each other. While these combined effects on bulk properties have received some attention [41, 42, 43], the effects on the micro-structure are not fully understood. In particular, studies of effects on cell characteristics in 2D have been limited so far to squares with rounded corners [44, 45].
A way to explore effects of particle angularity in 2D, is by observing packings of $N$-sided regular polygonal particles, studying the effects of modifying $N$, and checking consistency in the discs limit, $N\to\infty$. In particular, we investigate systematically the interplay between angularity and friction and which indicators of SO [33, 46, 3] survive the increasing angularity.
@@ II.1 Simulation setup
We performed biaxial compression tests on $N$-edged regular polygon systems, with $N=5,6,13$, and $64$. $N=5$ was chosen because it is the most angular, non-space filling shape. $N=64$ because the kinematics of such systems are hardly different from discs and $N=13$ was chosen to provide intermediate data between the former two. $N=6$ was chosen as a reference point because structures of granular media often display hexagonal symmetry in 2D. $N=3$ and $4$ were omitted because these readily organize into polycrystalline structures [47], sometimes also with bidisperse size distributions. We used the hard-particle soft-contact Discrete Element Method for our simulations [48, 33]. In each simulation, 2200-2750 frictionless particles of equal shape were deposited under gravity initially to generate dense packings. We used the same initial domain for all simulations, resulting in less angular particles starting from denser packings. To minimize crystallization, we used an 80%-20% small-to-large mixture of sizes in each run, with circum-radii $r=2.5$ and 55mm, respectively. Further simulation parameters are given in table 1. Once generated, gravity and sidewall friction were disabled to create isotropic conditions. The systems were compressed uniaxially by adjusting the positions of the left, right, and top walls until the pressure on all walls reached $P_{0}=20$kPa. Energy was then allowed to dissipate until pressure fluctuations decayed to less than $10^{-14}$ of the constant pressure. At this ’initial’ state, intergranular friction, $\mu$, was switched on and the packing was compressed biaxially by moving the top wall down at constant speed $v_{y}$ and adjusting the left and right wall to maintain constant $P_{0}$, as sketched in Fig. 1. $P_{0}$ was kept constant, using a servo-mechanism [49]. For each value of $N$, we simulated systems with five different friction coefficients, $\mu=0.01,0.1,0.5,1.0,10.0$, each starting from a different initial random configuration. In total, 20 different combinations of $N$ and $\mu$ were simulated, each five times, for good statistics. The pure shear, generated by the biaxial compression, continued well after the global stress plateaued out. The system’s height at time $t$, $h(t)$, decreases relative to $h(t=0)$ and the strain is defined as
$\epsilon=(h(t=0)-h(t))/h(t=0)\ .$ (1)
In all our runs, all measurable quantities reached a steady state by the time $\epsilon$ reached $8\%$. Therefore, all steady-state data were collected between strains of 88% and 1616%.
Table 1: Numerical parameters for the simulations.
Num. particles $\sim$ 2200-2750
Mixing ratio 80% small
20% large,
Circum-radii 2.5 & 5.0 [mm]
Corner numbers NN $[5,6,13,64]$
Friction coefficient $\mu$ $[0.01,0.10,0.50,$
$1.00,10.0]$
Young’s modulus $Y$ $10^{9}$ [N/m]
Density $\varrho$ $2830$ [kg/m2]
Confining pressure $P_{0}$ 2020 [kPa]
Compression velocity $v_{y}$ $10$ [mm/s]
Ratio $Y/P_{0}$ $5\cdot 10^{5}$
Timestep $\Delta t$ $2\cdot 10^{-6}$ [s]
Figure 1: The biaxial shear setup and boundary conditions. The top wall is lowered down at a constant speed $v_{y}$ and the side walls are free to move to maintain a constant confining pressure $P_{0}$.
@@ II.2 Cells and cell stresses
In DEM, particles can penetrate slightly into one another and a contact between two particles is defined as the centroid of their overlapping areas. A cell is then defined as the polygon whose vertices are the contacts between the particles surrounding it, e.g., the red-shaded area in Fig. 2. The cells in our simulations were generically convex, although this need not be the general case. The order of a cell, $k$, is defined as the number of particles surrounding it, $k\geq 3$. Our analysis is limited to $k\leq 8$ because of the relatively poorer statistics and short life of higher-order cells
We define a cell stress $\bm{\sigma}_{c}$ as follows. Starting from the stresses of the particles $p$ that surround cell $c$,
$\bm{\sigma}_{p}=\frac{1}{A_{p}}\sum_{p^{\prime}}\bm{l}_{pp^{\prime}}\otimes\bm{f}_{p^{\prime}p}\ ,$ (2)
with $p^{\prime}$ the particles in contact with $p$, $\bm{l}_{p^{\prime}p}$ vectors connecting the centroid of $p$ with the contact with $p^{\prime}$, and $\bm{f}_{p^{\prime}p}$ the force exerted by $p^{\prime}$ on $p$. The area associated with particle $p$, $A_{p}$, is the sum of the areas of its quadrons [50, 51]: $A_{pc},A_{pc^{\prime}},A_{pc"}$, which extend to its surrounding cells, respectively, $c,c^{\prime},c$", as shown in Fig. 2. In terms of these, the cell stress is
$\bm{\sigma}_{c}=\frac{\sum_{p\in c}A_{pc}\bm{\sigma}_{p}}{\sum_{p\in c}{A_{pc}}},$ (3)
with $p\in c$ denoting the particles surrounding cell $c$.
Figure 2: An example of a cell of order $k=8$: The cell is represented as a polygon, whose vertices are the centroids of the overlapping areas of its surrounding particles. $A_{pc},A_{pc^{\prime}}$, and $A_{pc"}$ (in red) are the areas of the quadrons of $p$ extending into its surrounding cells, respectively, $c,c^{\prime}$, and cc", and they represent the area associated with polygon pp. The cell area is the sum of the quadrons of the particles extending into it.
II.3 Visualization of cells and forces
In Fig. 3, we show in the left panel examples of cells for several combinations of $N$ and $\mu$. When $\mu=10.0$, angular particles readily form cells with $k>8$. The particles shown inside cells are rattlers, which are identified by visualizing the force chains network, outlined by the black lines, as illustrated in the right panel.
@@ III Results
We find that the conditional cell order probabilities, $P(k|N)$, depend on on both the angularity and on friction (Fig. 4) – the more angular the particles the more sensitive they are to $\mu$. Interestingly, while cells of order $3$ are the most frequent for $N\leq 13$, cells of order $4$ are the most frequent when $N$ increases. As expected, increasing $\mu$ also increases the occurrence probabilities of high-order cells, but up to cell order $5$, $P(k\leq 5|N)$ decreases with $\mu$ for all $N$. Reducing $N$ increases the proportion of order-$3$ cells and results in more compact structures.
@@
[…]
$P\left(\hat{h}_{k}\right)=\frac{m}{\lambda}\left(\frac{\hat{h}_{k}}{\lambda}\right)^{m-1}e^{-(\hat{h}_{k}/\lambda)^{m}}\ ,$ (4)
[…]
and found that these fits are worse than the Weibull one, with the parameters no longer independent of $N$ (Fig. 6).
Figure 5: The collapsed form of P(h^)P($\hat${h}) for different values of $\mu$ and $N$. The dashed lines represent Weibull-distribution fits for the different values of $\mu$.
Figure 6: NN-dependence of the parameters of the Weibull distribution fit to the cell probability distributions in Fig. 4 in the main text. For reference we also plot the parameters of a k-Gamma distribution to the same data. The error for the Weibull parameters is too small to be seen at this scale.
Fig. 7 illustrates the convergence of the cell conditional probability distributions to their steady-state values, $\mathrm{d}P(k|N,t)/\mathrm{d}t\to 0$, for different values of $N$ and $\mu$. The fluctuations around the steady state increase somewhat with $\mu$, which is expected to be a general feature. In contrast, the difference between the initial convergence of the very low-friction systems ($\mu=0.01$) and the other systems is attributed to the initial system preparation and the particular procedure we used to generate the steady state.
Figure 7: The rate of change of the cell conditional probabilities, $dP(k|N)/dt$, with strain for different combinations of friction $\mu$ and angularity NN.
In Fig. 8 we plot the mean cell stress ratio for each $k$, NN, and $\mu$. While the initial evolution of the cell stress-ratio, $\sigma_{c1}/\sigma_{c2}$, is qualitatively different for the different combinations of $N$ and $\mu$, the steady-state values were very similar. The exception from the above is the most angular particles assembly, $N=5$, which we discuss in detail below and in the concluding section. Nevertheless, the convergence in almost all cases to a very similar steady-state stress ratio, observed here for the first time, supports further the idea of cooperative stress-structure self-organisation [3].
@@
The steady-state values of the stress ratio $\sigma_{c1}/\sigma_{c2}$, shown in Fig. 9, depend non-monotonically on $\mu$. It increases with $\mu$ up to a maximum between $0.1\leq\mu\leq 0.5$ and decreases For $\mu\geq 1.0$. The rate of decrease depends on both cell order and particle shape. This is in line with earlier observations [52, 43] and can be understood in terms of a competition between inter-particle sliding and rolling which leads to an optimal stability around $\mu=0.5$. We elaborate on this mechanism below. This trend is less pronounced for $N=6$, where the stress ratio is also generally higher than for any other shape. This is because hexagons are more stable, which suppresses this competition.
Figure 9: Dependence of the steady state stresses ratio, shown in Fig. 8 on $\mu$ and $N$.
To investigate cell orientational organisation, we approximate each cell as the ellipse closest to its contacts [3] (Fig. 10). We measured the ellipses’ aspect ratios, $\alpha$, and major axes orientations, $\theta_{c}$. We omit in this analysis cells of order $k=3$ for which the ellipse approximation is not useful.
@@
Figure 11 shows the evolution of $\alpha$ with strain for combinations of $N$ and $\mu$. At low friction, $\alpha$ is constant regardless of $N$. As $\mu$ increases the mean aspect ratio increases and then decrease to a steady state value. The maximum is more pronounced for large $k$, low $N$, and high $\mu$. It is hardly observable or disappears altogether for $k\leq 5$. In this regime, $\alpha$ decreases with increasing cell order. This can be understood in the context of the entropy-mechanical stability competition [8], with the latter dominating at high-order cells.
@@
Figure 12: Cell aspect ratios at steady state as a function of $\mu$ for different angularity NN.
Cell organisation is driven by stresses and we find that, for all $N$ and $\mu$, there is a strong correlation between cell orientations $\theta_{c}$ and the local major stress principal axis $\theta_{s}$ (see Fig. 13). As Fig. 14 shows, the correlation becomes weaker as $\mu$ decreases. This is both because cells are less elongated and because the media around cells are more mobile and allow easier cell reorientation.
@@
In Fig. 14 we plot the distribution of the local difference between the two orientations, $\Delta\theta\equiv\theta_{c}-\theta_{s}$, establishing beyond doubt their tendency to align. This phenomenon, which is more pronounced at high-order cells, is as in disc systems [2, 3], suggesting that it is independent of particle angularity.
@@
We understand the dynamics that lead to such organisation as follows: Although the process is quasistatic, reorientation of cells requires unbalanced forces to move particles. This means that the cell stresses are momentarily asymmetric, giving rise to small torques that rotate them. The less stable the cell the larger the rotation. To quantify a cell’s stability, let the axis of its $\sigma_{c1}$ be at an angle $\Delta\theta$ to its major ellipse axis. In the frame of reference aligned with the latter, the cell stress is
$\bm{\Sigma}=\begin{bmatrix}\sigma_{c1}+\sigma_{c2}\tan^{2}\Delta\theta&2q\tan\Delta\theta\\ 2q\tan\Delta\theta&\sigma_{c1}\tan^{2}\Delta\theta+\sigma_{c2}\end{bmatrix}\cos^{2}\Delta\theta\ .$ (6)
A common measure of its stability is the ratio of the off-diagonal stress component to the trace, in this case, $\tau=\left(\sigma_{c1}-\sigma_{c2}\right)\sin{\left(2\Delta\theta\right)}/\left[2\left(\sigma_{c1}+\sigma_{c2}\right)\right]$. The lower $\tau$ is the more stable the cell. During the dynamic process, the cell is nudged towards higher stability, which amounts to reducing $\Delta\theta$. It is this nudging that gives rise to the observed peak around $\Delta\theta=0$. This can also be seen in the scatter plot of cells in the plane spanned by $\Delta\theta$ and $\tau/p$, shown in Fig. 15.
Figure 14: PDF of $\Delta\theta\equiv\theta_{c}-\theta_{s}$ for different cell orders, $k$.
@@
Figure 15: A scatter plot heat map of cells in the plane spanned by $\Delta\theta$ and $\tau/p$ plane. The colour denotes the relative frequency of cells with given values of $\Delta\theta$ and $\tau/p$.
Generally, there is a competition between two cell rearrangement mechanism [43]. Particles contacting along edges can slide relative to one another while corner-edge and (the rarer) corner-corner contacts can change mainly by a relative rotation. For a given friction coefficient, the more angular the particles the easier it is to rearrange by rotation. Increasing $\mu$ for systems of the same $N$, would also favour rotation over sliding. However, as $\mu\to 0$ for a given $N$, the rotation-vs-sliding competition goes from being determined by the two particles in contact to depending on the complex many-body dynamics of their environment, which becomes more fluid. It is the interplay between these mechanisms that leads to the difference between the behaviour of the $N=5$ particles and that of high-$N$ and discs.